# Cyclotomic generating functions

### v2

@@
Abstract.
It is a remarkable fact that for many statistics on finite sets of combinatorial objects, the roots of the corresponding generating function are each either a complex root of unity or zero. These and related polynomials have been studied for many years by a variety of authors from the fields of combinatorics, representation theory, probability, number theory, and commutative algebra. We call such polynomials cyclotomic generating functions (CGF’s). In [BKS20a, BKS20b, BS22], the authors studied the support and asymptotic distribution of the coefficients of several families of CGF’s arising from tableau and forest combinatorics. In this paper, we survey general CGF’s from algebraic, analytic, and asymptotic perspectives. We review some of the many known examples of CGF’s in combinatorial representation theory; describe their coefficients, moments, cumulants, and characteristic functions; and give a variety of necessary and sufficient conditions for their existence arising from probability, commutative algebra, and invariant theory. As a sample result, we show that CGF’s are “generically” asymptotically normal, generalizing a result of Diaconis on $q$-binomial coefficients using work of Hwang–Zacharovas. We include several open problems concerning CGF’s.
Key words and phrases:
cyclotomic polynomials, $q$-integers, asymptotic normality, cumulants
The authors were partially supported by the Washington Research Foundation and DMS-1764012.
1. Introduction
Many formulas in enumerative combinatorics express the cardinality of a finite set $X$ as a product or quotient of integers. In many cases of interest, such formulas may be generalized to $q$-analogues with a corresponding refined count of $X$ subject to the value of some statistic on $X$. Such $q$-analogues have non-negative integer coefficients. A classic example is the $q$-binomial coefficients, which $q$-count integer partitions that fit in the $k\times(n-k)$ rectangle according to size. Here a partition $\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{j}\geq 0)$ fits in the rectangle provided $j\leq k$ and each $\lambda_{i}$ is a positive integer less than or equal to $(n-k)$. The size of $\lambda$ is the sum of its parts, $|\lambda|=\sum\lambda_{i}$. These polynomials can be expressed in two ways,
(1.1) $\binom{n}{k}_{q}\coloneqq\frac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}=\sum_{\lambda\subset k\times(n-k)}q^{|\lambda|},$
@@ 1. Introduction
In this paper, we study the class of all such nonzero polynomials with non-negative integer coefficients arising as quotients of $q$-integers from algebraic, analytic, and asymptotic perspectives. Such polynomials are closely associated with the cyclotomic polynomials $\Phi_{n}$ from number theory. See Section 2 for a review of their key properties. Since our examples all come from $q$-counting formulas for well-known combinatorial objects and their associated generating functions, we have chosen to call these polynomials cyclotomic generating functions or CGF’s for short. We begin by stating several equivalent characterizations of this class. See Section 3 for the proof of their equivalence.
Theorem/Definition 1.1.
Suppose $f(q)=\sum_{k=0}^{n}c_{k}q^{k}$ is a nonzero polynomial with non-negative integer coefficients. The following are equivalent definitions for the polynomial $f$ to be a cyclotomic generating function (CGF).
@@
    (Rational form.) There are multisets $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{m}\}$ of positive integers and integers $\alpha\in\mathbb{Z}_{>0}$, $\beta\in\mathbb{Z}_{\geq 0}$ such that
@@
    where the summands are all independent and $\beta$ is maximal such that $q^{\beta}\mid f(q)$. Here $\mathcal{U}_{a}$ is a uniform random variable supported on $\{0,\ldots,a-1\}$ and $\alpha$ is the greatest common divisor of all coefficients of $f$.
@@
    (Characteristic form.) There is a discrete random variable $\mathcal{X}$ on a uniform sample space supported on $\mathbb{Z}_{\geq 0}$ with probability generating function $\mathbb{E}[q^{\mathcal{X}}]=f(q)/f(1)$ such that the scaled characteristic function $\phi_{\mathcal{X}}(z):=\mathbb{E}[e^{2\pi iz\mathcal{X}}]$ has complex zeros only at rational $z\in\mathbb{Q}$.
Example 1.2.
Allow us to indulge in a thought experiment. Imagine yourself as Percy MacMahon at the turn of the last century investigating plane partitions in an $x\times y\times z$ box for the first time. These are $x\times y$ matrices of non-negative integers whose entries are at most $z$ and which weakly decrease along rows and columns. For $(x,y,z)=(3,2,2)$, you find there are $50$ plane partitions. We have $50=2\cdot 5^{2}$, and a wide-eyed optimist may hope for a product formula akin to that for $\binom{n}{k}$, though cancellations are difficult to uncover. Plane partitions come with the size statistic, which is the sum of all entries. The size generating function here is
@@
(1.2) $f_{(x,y,z)}(q)=\prod_{i=1}^{x}\prod_{j=1}^{y}\frac{[i+j+z-1]_{q}}{[i+j-1]_{q}},$
@@
Definition 1.3.
Among all CGF’s, we focus on the family of basic CGF’s with no $\alpha$ or $q^{\beta}$ factors,
(1.3) $f(q)=\prod_{j=1}^{m}\frac{[a_{j}]_{q}}{[b_{j}]_{q}}=\prod_{j=1}^{m}\frac{1-q^{a_{j}}}{1-q^{b_{j}}}.$
The basic CGF monoid $\Phi^{+}$ is the monoid consisting of all basic CGF’s under multiplication.
By the rational form, basic CGF’s are always palindromic and monic, though they need not be unimodal in general, e.g.
(1.4) $f(q)=q^{6}+q^{4}+q^{3}+q^{2}+1=\frac{[5]_{q}[6]_{q}}{[2]_{q}[3]_{q}}\in\Phi^{+}.$
Testing if a polynomial $f\in\mathbb{Z}_{\geq 0}[q]$ is in $\Phi^{+}$ may be done by repeatedly dividing off cyclotomic polynomial factors $\Phi_{n}(q)$. Note that we have the elementary bounds $\sqrt{n/2}\leq\deg\Phi_{n}(q)\leq n$, so there are only a finite number of basic CGF’s of each given degree.
The main reason to study CGF’s as a family of polynomials is that they are already prevalent in the literature. In addition to $q$-binomial coefficients, the standard $q$-analogues of $n!$ using inversions or the major index statistic on permutations and their generalizations to arbitrary words corresponding with $q$-multinomial generating functions are CGF’s using MacMahon’s classic formulas. Here are a Catalan number of examples from the literature. We emphasize that this list is not exhaustive.
  1. (1)
@@
  2. (2)
    The Hilbert series of all finite-dimensional quotient rings of the form $R\coloneqq B/(\theta_{1},\ldots,\theta_{m})$ where $\theta_{1},\ldots,\theta_{m}$ is a homogeneous system of parameters in the polynomial ring $B=\mathbbm{k}[x_{1},\ldots,x_{m}]$, where $\deg(\theta_{i})=a_{i}$ and $\deg(x_{i})=b_{i}$. The ring of invariants of a finite reflection group has such a Hilbert series. See Section 6.3 for references and further discussion.
  3. (3)
@@
  14. (14)
    Rank generating functions of Gaussian posets and $d$-complete posets, with connections to Lie theory and order polynomials of posets, [Pro84, PS19, Ste94]. In recent work, Hopkins has explored general properties of posets with order polynomial product formulas in the context of cyclic sieving and other good dynamical behavior such as promotion and rowmotion [Hop23]. Furthermore, sometimes there is an associated cyclic sieving phenomena associated to these order polynomials. These well behaved order polynomials are often cyclotomic generating functions. See also Stanton’s online notes [Sta98] for conjectures related to CGF’s generalizing rank generating functions of Gaussian posets, which he calls “fake Gaussian sequences”.
Classes of polynomials similar to the class of cyclotomic generating functions have been studied for roughly a century in other contexts. A polynomial $f\in\mathbb{C}[q]$ of degree $n$ is self-inversive if $f(q)=\alpha q^{n}f(q^{-1})$ where $|\alpha|=1$. Equivalently, the zeros of $f$ are symmetric about the unit circle. Self-inversive polynomials were studied by A. Cohn [Coh22] a century ago and by Bonsall–Marden [BM52] in the 1950s; they related them to the complex roots of $f^{\prime}$. Kedlaya [Ked08] and Hwang–Zacharovas [HZ15] refer to polynomials with roots all on the unit circle as root-unitary. Note that basic CGF’s are in particular root-unitary with real coefficients. We will review key details from [HZ15] in Section 2 since they pertain to our focus on basic CGF’s. In particular, they gave an elegant formula for the cumulants of the real-valued discrete random variables whose probability generating functions are basic CGF’s and a test for asymptotic normality using the fourth cumulants. More recently, there has been interest in certain polynomials associated to numerical semigroups that have cyclotomic factorizations [CGSM16, CGSHPM22] and identifying criteria for central limit theorems for polynomials based on the geometry of their roots in the complex plane [HST23, MS19b, MS19a].
Another particularly well-known class of polynomials consists of elements of $\mathbb{R}_{\geq 0}[q]$ with all real roots [Brä15, Bre94, Pit97, Sta89]. Stirling numbers of the second kind are the motivating example [Har67]. A famous asymptotic characterization of their coefficients is due to Bender.
Theorem 1.4 ([Ben73, Thm. 2]; see also [Har67]).
@@
In contrast to 1.4, CGF polynomials and their associated random variables have much more complex limiting behavior. In [BS22], we initiated the study of the metric space of all standardized CGF distributions in the Lévy metric and their asymptotic limits. For some families of CGF’s, there are statistics that completely characterize their limiting behavior. Asymptotic normality is one common occurrence. While finding a complete description of the closure of all CGF distributions is still an open problem [BS22, Open Problem 1.19], we did show that all uncountably many DUSTPAN distributions can occur on the boundary. Furthermore, many more multimodal limiting distributions are possible, due to the following construction. Given a CGF $f(q)$ and positive integer $N$, note that $(1+q^{N})f(q)$ remains a CGF. If $N$ is larger than the degree of $f(q)$, the result is a CGF with two disjoint copies of the distribution of $f(q)$. In this way, valid limits may be “duplicated” in a fractal pattern. The unimodal and log-concave submonoids of $\Phi^{+}$ are not subject to this construction. We expect the unimodal basic CGF’s will have the simplest limiting behavior.
Towards addressing the problem of characterizing when a limiting sequence of standardized CGF distributions in the Lévy metric approaches the normal distribution $\mathcal{N}(0,1)$, we present a new criterion in terms of the multisets in the rational presentation of the CGF polynomials. It is these multisets which have combinatorial significance in applications. We give a more complete characterization in the special case of polynomials which are products of $q$-integers, see 4.1. This approach builds on the work of Hwang–Zacharovas [HZ15], but emphasizes the role of the multisets of $q$-integers in the numerators. It also generalizes the work of Diaconis [Dia88, pp.128-129] who showed that the coefficients of a sequence of $q$-binomials $\binom{n}{k}_{q}$ are asymptotically normal provided both $k,n-k\to\infty$. See Section 4 on asymptotic considerations for the proof.
Theorem 1.5.
@@
(1.5) $\limsup_{N\to\infty}\frac{\sum_{b\in b^{(N)}}(b^{2}-1)}{\sum_{a\in a^{(N)}}(a^{2}-1)}<1$
@@
(1.6) $\lim_{N\to\infty}\sum_{a\in a^{(N)}}\left(\frac{a}{\max a^{(N)}}\right)^{4}=\infty,$
@@
When $x\leq y\leq z$, which may be arranged without loss of generality, and $z\to\infty$, we use the rational function above to observe (1.5) holds. Similarly we find that (1.6) holds if $y\to\infty$ as well. Hence with nothing more than basic computer algebra systems and routine calculations, we find that size on plane partitions in a box is asymptotically normal if $\mathrm{median}(x,y,z)\to\infty$, recovering a result from [BS22]. Indeed, this result is sharp in the sense that, when $\mathrm{median}(x,y,z)$ is bounded and $z\to\infty$, the possible limits are Irwin–Hall distributions; see [BS22]. The approach in this example to proving central limit theorems is a kindred spirit to Zeilberger’s “automatic central limit theorem generator” [Zei09].
Remark 1.7.
While 1.5 provides a precise asymptotic conclusion, [MS19a, Thm. 1.4] states the following quantitative bound after a lengthy technical argument. If $f(q)$ is a polynomial which is zero-free in the sector $\{z\in\mathbb{C}:|\arg(q)|<\delta\}$ and if $\mathcal{X}$ is a random variable with $\mathbb{E}[q^{\mathcal{X}}]=f(q)/f(1)$, then
(1.7) $\sup_{t\in\mathbb{R}}|\mathbb{P}(\mathcal{X}^{*}\leq t)-\mathbb{P}(\mathcal{Z}\leq t)|\leq\frac{C}{\delta\sigma},$
where $C$ is an absolute constant ant $\mathcal{Z}$ is the standard normal. In the situation of 1.5, $\delta^{-1}=\max a^{(N)}/2\pi$ and $12\sigma^{2}=\sum_{a\in a^{(N)}}a^{2}-\sum_{b\in b^{(N)}}b^{2}$, so (1.7) becomes
(1.8) $\sup_{t\in\mathbb{R}}|\mathbb{P}(\mathcal{X}^{*}\leq t)-\mathbb{P}(\mathbb{Z}\leq t)|\leq\frac{C_{0}\max a^{(N)}}{\sqrt{\sum_{a\in a^{(N)}}a^{2}-\sum_{b\in b^{(N)}}b^{2}}}.$
@@
Theorem 1.8.
Let $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ be a sequence of random variables corresponding to cyclotomic generating functions. Then the sequence of standardized random variables $\mathcal{X}_{1}^{*},\mathcal{X}_{2}^{*},\ldots$ converges in distribution if and only if for all $d\in\mathbb{Z}_{\geq 1}$,
@@
exists and is finite. In this case, $\mathcal{X}_{n}^{*}\Rightarrow\mathcal{X}$ where $\mu_{d}=\mu_{d}^{\mathcal{X}}$, $\mathcal{X}$ is determined by its moments, and the moment-generating function of $\mathcal{X}$ is entire.
The paper is organized as follows. In Section 2, we state our notation and give background details from prior work. In Sections 3,  4, and 5, we describe the algebraic, asymptotic and analytic properties of basic CGF’s, their associated random variables, and characteristic functions. In Section 6, we examine the monoid of basic cyclotomic generating functions and several of its submonoids including unimodal and log-concave CGF’s, CGF’s whose rational form satisfies the Gale order, and CGF’s that come from Hilbert series of polynomial rings modded out by homogeneous systems of parameters. We include several open problems about cyclotomic generating functions for future work based on experimentation.
2. Background
2.1. Cyclotomic polynomials
The cyclotomic polynomials are defined for all positive integers $n\geq 1$ by
(2.1) $\Phi_{n}(q)=\prod_{\begin{subarray}{c}1\leq k\leq n\\ \gcd(k,n)=1\end{subarray}}(q-e^{2\pi ik/n}).$
Each $n^{th}$ root of unity $e^{2\pi ik/n}$ is in the cyclic group generated by $e^{2\pi i/d}$ for $d\in[n]$ where $d=n/\gcd(k,n)$. Thus, for $n\geq 1$, we have
(2.2) $q^{n}-1=\prod_{k=1}^{n}(q-e^{2\pi ik/n})=\prod_{d|n}\Phi_{d}(q).$

### v3

@@
Abstract
It is a remarkable fact that for many statistics on finite sets of combinatorial objects, the roots of the corresponding generating function are each either a complex root of unity or zero. These and related polynomials have been studied for many years by a variety of authors from the fields of combinatorics, representation theory, probability, number theory, and commutative algebra. We call such polynomials cyclotomic generating functions (CGFs). With Konvalinka, we have studied the support and asymptotic distribution of the coefficients of several families of CGFs arising from tableau and forest combinatorics. In this paper, we survey general CGFs from algebraic, analytic, and asymptotic perspectives. We review some of the many known examples of CGFs in combinatorial representation theory; describe their coefficients, moments, cumulants, and characteristic functions; and give a variety of necessary and sufficient conditions for their existence arising from probability, commutative algebra, and invariant theory. As a sample result, we show that CGFs are “generically” asymptotically normal, generalizing a result of Diaconis on $q$-binomial coefficients using work of Hwang–Zacharovas. We include several open problems concerning CGFs.
1 Introduction
Many formulas in enumerative combinatorics express the cardinality of a finite set $X$ as a product or quotient of integers. In many cases of interest, such formulas may be generalized to $q$-analogues with a corresponding refined count of $X$ subject to the value of some statistic on $X$. Such $q$-analogues have nonnegative integer coefficients. A classic example is the $q$-binomial coefficients, which $q$-count integer partitions that fit in the $k\times(n-k)$ rectangle according to size for a given pair of nonnegative integers $k\leqslant n$. Here a partition $\lambda=(\lambda_{1}\geqslant\cdots\geqslant\lambda_{j}\geqslant 0)$ fits in the rectangle provided $j\leqslant k$ and each $\lambda_{i}$ is a nonnegative integer less than or equal to $(n-k)$. The size of $\lambda$ is the sum of its parts, $|\lambda|=\sum\lambda_{i}$. These polynomials can be expressed in two ways,
$\binom{n}{k}_{q}\coloneqq\frac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}=\sum_{\lambda\subset k\times(n-k)}q^{|\lambda|},$ (1.1)
@@ 1 Introduction
In this paper, we study the class of such nonzero polynomials with nonnegative integer coefficients arising as quotients of $q$-integers from algebraic, analytic, and asymptotic perspectives. Such polynomials are closely associated with the cyclotomic polynomials $\Phi_{n}$ from number theory. See Section 2 for a review of their key properties. Since our examples all come from $q$-counting formulas for well-known combinatorial objects and their associated generating functions, we have chosen to call these polynomials cyclotomic generating functions or CGFs for short. We begin by stating several equivalent characterizations of this class. See Section 3 for the proof of their equivalence.
Theorem/Definition 1.
Suppose $f(q)=\sum_{k=0}^{n}c_{k}q^{k}$ is a nonzero polynomial with nonnegative integer coefficients. The following are equivalent definitions for the polynomial $f$ to be a cyclotomic generating function (CGF).
@@
    (Rational form.) There are multisets $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{m}\}$ of positive integers and integers $\alpha\in\mathbb{Z}_{>0}$, $\beta\in\mathbb{Z}_{\geqslant 0}$ such that
@@
    where the summands are all independent and $\beta$ is maximal such that $q^{\beta}\mid f(q)$. Here $\mathcal{U}_{a}$ is a uniform random variable supported on $\{0,1,\ldots,a-1\}$ and $\alpha$ is the greatest common divisor of all coefficients of $f$.
@@
    (Characteristic form.) There is a discrete random variable $\mathcal{X}$ on a uniform sample space supported on $\mathbb{Z}_{\geqslant 0}$ with probability generating function $\mathbb{E}[q^{\mathcal{X}}]=f(q)/f(1)$ such that the scaled characteristic function $\phi_{\mathcal{X}}(z):=\mathbb{E}[e^{2\pi iz\mathcal{X}}]$ has complex zeros only at rational $z\in\mathbb{Q}$.
Example 2.
Allow us to indulge in a thought experiment. Imagine yourself as Percy MacMahon at the turn of the last century investigating plane partitions in an $x\times y\times z$ box for the first time. These are $x\times y$ matrices of nonnegative integers whose entries are at most $z$ and weakly decrease along rows and columns. For $(x,y,z)=(3,2,2)$, you find there are $50$ plane partitions. We have $50=2\cdot 5^{2}$, and a wide-eyed optimist may hope for a product formula akin to that for $\binom{n}{k}$, though cancellations are difficult to uncover. Plane partitions come with the size statistic, which is the sum of all entries. The size generating function here is
@@
$f_{(x,y,z)}(q)=\prod_{i=1}^{x}\prod_{j=1}^{y}\frac{[i+j+z-1]_{q}}{[i+j-1]_{q}},$ (1.2)
@@
Definition 3.
Among all CGFs, we focus on the family of basic CGFs with no $\alpha$ or $q^{\beta}$ factors,
$f(q)=\prod_{j=1}^{m}\frac{[a_{j}]_{q}}{[b_{j}]_{q}}=\prod_{j=1}^{m}\frac{1-q^{a_{j}}}{1-q^{b_{j}}}.$ (1.3)
The basic CGF monoid $\Phi^{+}$ is the monoid consisting of all basic CGFs under multiplication.
By the rational form, basic CGFs are always palindromic and monic. They need not be unimodal in general, e.g.
$f(q)=q^{6}+q^{4}+q^{3}+q^{2}+1=\frac{[5]_{q}[6]_{q}}{[2]_{q}[3]_{q}}\in\Phi^{+}.$ (1.4)
In Section 6.1, we study the submonoids $\Phi^{\operatorname{uni}}$ and $\Phi^{\operatorname{lcc}}$ of $\Phi^{+}$ given by cyclotomic generating functions which are unimodal or log-concave with no internal zeros, respectively.
Testing if a polynomial $f\in\mathbb{Z}_{\geqslant 0}[q]$ is in $\Phi^{+}$ may be done by repeatedly dividing off cyclotomic polynomial factors $\Phi_{n}(q)$. Note that we have the elementary bounds $\sqrt{n/2}\leqslant\deg\Phi_{n}(q)\leqslant n$, so there are only a finite number of basic CGFs of each given degree.
The main reason to study CGFs as a family of polynomials is that they are already prevalent in the literature. In addition to $q$-binomial coefficients, the standard $q$-analogues of $n!$ using inversions or the major index statistic on permutations and their generalizations to arbitrary words corresponding to $q$-multinomial generating functions are CGFs using MacMahon’s classic formulas. Here are a (Catalan) number of examples from the literature. We emphasize that this list is not exhaustive.
  1. 1.
@@
  2. 2.
    The Hilbert series of all finite-dimensional quotient rings of the form
    $R\coloneqq B/(\theta_{1},\ldots,\theta_{m})$
    where $\theta_{1},\ldots,\theta_{m}$ is a homogeneous system of parameters in the polynomial ring $B=\mathbbm{k}[x_{1},\ldots,x_{m}]$, where $\deg(\theta_{i})=a_{i}$ and $\deg(x_{i})=b_{i}$. The ring of invariants of a finite reflection group has such a Hilbert series. See Section 6.3 for references and further discussion.
  3. 3.
@@
  14. 14.
    Rank generating functions of Gaussian posets and $d$-complete posets, with connections to Lie theory and order polynomials of posets, [Pro84, PS19, Ste94]. In recent work, Hopkins has explored general properties of posets with order polynomial product formulas in the context of cyclic sieving and other good dynamical behavior such as promotion and rowmotion [Hop24]. Furthermore, sometimes there is an associated cyclic sieving phenomenon associated to these order polynomials. These well behaved order polynomials are often cyclotomic generating functions. See also Stanton’s online notes [Sta98] for conjectures related to CGFs generalizing rank generating functions of Gaussian posets, which he calls “fake Gaussian sequences”.
Classes of polynomials similar to the class of cyclotomic generating functions have been studied for roughly a century in other contexts. A polynomial $f\in\mathbb{C}[q]$ of degree $n$ is self-inversive if $f(q)=\alpha q^{n}f(q^{-1})$ where $|\alpha|=1$. Equivalently, the zeros of $f$ are symmetric about the unit circle. Self-inversive polynomials were studied by A. Cohn [Coh22] a century ago and by Bonsall–Marden [BM52] in the 1950s; they related them to the complex roots of $f^{\prime}$.
Kedlaya [Ked08] and Hwang–Zacharovas [HZ15] refer to polynomials with roots all on the unit circle as root-unitary. Note that basic CGFs are in particular root-unitary with real coefficients. We will review key details from [HZ15] in Section 2 since they pertain to our focus on basic CGFs. In particular, they gave an elegant formula for the cumulants of the real-valued discrete random variables whose probability generating functions are basic CGFs and a test for asymptotic normality using the fourth cumulants. More recently, there has been interest in certain polynomials associated to numerical semigroups that have cyclotomic factorizations [CGSM16, CGSHPM22] and identifying criteria for central limit theorems for polynomials based on the geometry of their roots in the complex plane [HST24, MS19b, MS19a].
Another particularly well-known class of polynomials consists of elements of $\mathbb{R}_{\geqslant 0}[q]$ with all real roots [Brä15, Bre94, Pit97, Sta89]. Stirling numbers of the second kind are the motivating example [Har67]. A famous asymptotic characterization of their coefficients is due to Bender.
Theorem 4 ([Ben73, Thm. 2]; see also [Har67]).
@@
In contrast to Theorem 4, CGF polynomials and their associated random variables have much more complex limiting behavior. In [BS22], we initiated the study of the metric space of all standardized CGF distributions in the Lévy metric and their asymptotic limits. For some families of CGFs, there are statistics that completely characterize their limiting behavior. Asymptotic normality is one common occurrence. While finding a complete description of the closure of all CGF distributions is still an open problem [BS22, Open Problem 1.19], we did show that all uncountably many DUSTPAN distributions (distributions associated to a uniform sum for $\mathbf{t}$ plus a normal distribution) can occur on the boundary. Furthermore, many more multimodal limiting distributions are possible, due to the following construction. Given a CGF $f(q)$ and positive integer $N$, note that $(1+q^{N})f(q)$ remains a CGF. If $N$ is larger than the degree of $f(q)$, the result is a CGF with two disjoint copies of the distribution of $f(q)$. In this way, valid limits may be “duplicated” in a fractal pattern. The unimodal and log-concave submonoids of $\Phi^{+}$ are not subject to this construction. We expect the unimodal basic CGFs will have the simplest limiting behavior.
Towards addressing the problem of characterizing when a limiting sequence of standardized CGF distributions in the Lévy metric approaches the standard normal distribution $\mathcal{N}(0,1)$, we present a new criterion in terms of the multisets in the rational presentation of the CGF polynomials. It is these multisets which have combinatorial significance in applications. We give a more complete characterization in the special case of polynomials which are products of $q$-integers, see Theorem 37. This approach builds on the work of Hwang–Zacharovas [HZ15], but emphasizes the role of the multisets of $q$-integers in the numerators. It also generalizes the work of Diaconis [Dia88, pp.128-129] who showed that the coefficients of a sequence of $q$-binomials $\binom{n}{k}_{q}$ are asymptotically normal provided both $k,n-k\to\infty$. See Section 4 on asymptotic considerations for the proof.
Theorem 5.
@@
$\limsup_{N\to\infty}\frac{\sum_{b\in b^{(N)}}(b^{2}-1)}{\sum_{a\in a^{(N)}}(a^{2}-1)}<1$ (1.5)
@@
$\lim_{N\to\infty}\sum_{a\in a^{(N)}}\left(\frac{a}{\max a^{(N)}}\right)^{4}=\infty,$ (1.6)
@@
When $x\leqslant y\leqslant z$, which may be arranged without loss of generality, and $z\to\infty$, we use the rational function above to observe (1.5) holds. Similarly we find that (1.6) holds if $y\to\infty$ as well. Hence with nothing more than basic computer algebra systems and routine calculations, we find that size on plane partitions in a box is asymptotically normal if $\mathrm{median}(x,y,z)\to\infty$, recovering a result from [BS22]. Indeed, this result is sharp in the sense that, when $\mathrm{median}(x,y,z)$ is bounded and $z\to\infty$, the possible limits are Irwin–Hall distributions; see [BS22]. The approach in this example to proving central limit theorems is a kindred spirit to Zeilberger’s “automatic central limit theorem generator” [Zei09].
Remark 7.
While Theorem 5 provides a precise asymptotic conclusion, [MS19a, Thm. 1.4] states the following quantitative bound after a lengthy technical argument. If $f(q)$ is a polynomial which is zero-free in the sector $\{z\in\mathbb{C}:|\arg(q)|<\delta\}$ and if $\mathcal{X}$ is a random variable with $\mathbb{E}[q^{\mathcal{X}}]=f(q)/f(1)$, then
$\sup_{t\in\mathbb{R}}|\mathbb{P}(\mathcal{X}^{*}\leqslant t)-\mathbb{P}(\mathcal{Z}\leqslant t)|\leqslant\frac{C}{\delta\sigma},$ (1.7)
where $C$ is an absolute constant and $\mathcal{Z}$ is the standard normal random variable. In the situation of Theorem 5, $\delta^{-1}=\max a^{(N)}/2\pi$ and $12\sigma^{2}=\sum_{a\in a^{(N)}}a^{2}-\sum_{b\in b^{(N)}}b^{2}$, so (1.7) becomes
$\sup_{t\in\mathbb{R}}|\mathbb{P}(\mathcal{X}^{*}\leqslant t)-\mathbb{P}(\mathbb{Z}\leqslant t)|\leqslant\frac{C_{0}\max a^{(N)}}{\sqrt{\sum_{a\in a^{(N)}}a^{2}-\sum_{b\in b^{(N)}}b^{2}}}.$ (1.8)
@@
Theorem 8.
Let $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ be a sequence of random variables corresponding to cyclotomic generating functions. Then the sequence of standardized random variables $\mathcal{X}_{1}^{*},\mathcal{X}_{2}^{*},\ldots$ converges in distribution if and only if for all $d\in\mathbb{Z}_{\geqslant 1}$, the limit of the $d^{th}$ moments of the random variables
@@
exists and is finite. In this case, $\mathcal{X}_{n}^{*}\Rightarrow\mathcal{X}$ where $\mu_{d}=\mu_{d}^{\mathcal{X}}$ is the $d^{th}$ moment of $\mathcal{X}$, $\mathcal{X}$ is determined by its moments, and the moment-generating function of $\mathcal{X}$ is entire.
The paper is organized as follows. In Section 2, we state our notation and give background details from prior work. In Sections 3, 4 and 5, we describe the algebraic, asymptotic and analytic properties of basic CGFs, their associated random variables, and characteristic functions. In Section 6, we examine the monoid of basic cyclotomic generating functions and several of its submonoids including unimodal and log-concave CGFs, CGFs whose rational form satisfies the Gale order, and CGFs that come from Hilbert series of polynomial rings modded out by homogeneous systems of parameters. We include several open problems about cyclotomic generating functions for future work based on experimentation.
2 Background
2.1 Cyclotomic polynomials
The cyclotomic polynomials are defined for all positive integers $n\geqslant 1$ by
$\Phi_{n}(q)=\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ \gcd(k,n)=1\end{subarray}}(q-e^{2\pi ik/n}).$ (2.1)
Each $n^{th}$ root of unity $e^{2\pi ik/n}$ is in the cyclic group generated by $e^{2\pi i/d}$ for $d\in\{1,2,\dots,n\}$ where $d=n/\gcd(k,n)$. Thus, for $n\geqslant 1$, we have
$q^{n}-1=\prod_{k=1}^{n}(q-e^{2\pi ik/n})=\prod_{d|n}\Phi_{d}(q).$ (2.2)